### v2

# Style-NeRF2NeRF: 3D Style Transfer from Style-Aligned Multi-View Images

Abstract
We propose a simple yet effective pipeline for stylizing a 3D scene, harnessing the power of 2D image diffusion models. Given a NeRF model reconstructed from a set of multi-view images, we perform 3D style transfer by refining the source NeRF model using stylized images generated by a style-aligned image-to-image diffusion model. Given a target style prompt, we first generate perceptually similar multi-view images by leveraging a depth-conditioned diffusion model with an attention-sharing mechanism. Next, based on the stylized multi-view images, we propose to guide the style transfer process with the sliced Wasserstein loss based on the feature maps extracted from a pre-trained CNN model. Our pipeline consists of decoupled steps, allowing users to test various prompt ideas and preview the stylized 3D result before proceeding to the NeRF fine-tuning stage. We demonstrate that our method can transfer diverse artistic styles to real-world 3D scenes with competitive quality. Result videos are also available on our project page: https://haruolabs.github.io/style-n2n/
Keywords Neural Radiance Fields  $\cdot$ Style Transfer  $\cdot$ Sliced Wasserstein
1 Introduction
Thanks to recent advancements in 3D reconstruction techniques such as Neural Radiance Fields (NeRF) Mildenhall et al. (2020), it is nowadays possible for creators to develop a 3D asset or a scene from captured real-world data without intensive labor. While such 3D reconstruction methods work well, editing an entire 3D scene to match a desired style or concept is not straightforward.
For instance, editing conventional 3D scenes based on explicit representations like mesh often involves specialized tools and skills. Changing the appearance of the entire mesh-based scene would often require skilled labor, such as shape modeling, texture creation, and material parameter modifications.
At the advent of implicit 3D representation techniques such as NeRF, style editing methods for 3D are also emerging Nguyen-Phuoc et al. (2022); Wang et al. (2023); Liu et al. (2023); Kamata et al. (2023); Haque et al. (2023); Dong and Wang (2024) to enhance creators’ content development process. Following the recent development of 2D image generation models, prominent works such as Instruct-NeRF2NeRF Haque et al. (2023); Vachha and Haque (2024) and ViCA-NeRF Dong and Wang (2024) proposed to leverage the knowledge of large-scale pre-trained text-to-image (T2I) models to supervise the 3D NeRF editing process.
These methods employ a custom pipeline based on an instruction-based T2I model ”Instruct-Pix2Pix” Brooks et al. (2023) to stylize a 3D scene with text instructions. While Instruct-NeRF2NeRF is proven to work well for editing 3D scenes including large-scale 360 environments, their method involves an iterative process of editing and replacing the training data during NeRF optimization, occasionally resulting in unpredictable results. As editing by Instruct-Pix2Pix runs in tandem with NeRF training, we found adjusting or testing editing styles beforehand difficult.
To overcome this problem, we propose an artistic style-transfer method that trains a source 3D NeRF scene on stylized images prepared in advance by a text-guided style-aligned diffusion model. Training is guided by Sliced Wasserstein Distance (SWD) loss Heitz et al. (2021); Li et al. (2022) to effectively perform 3D style transfer with NeRF. A summary of our contributions is as the follows:
  • We propose a novel 3D style-transfer approach for NeRF, including large-scale outdoor scenes.
  • We show that a style-aligned diffusion model conditioned on depth maps of corresponding source views can generate perceptually view-consistent style images for fine-tuning the source NeRF. Users can test stylization ideas with the diffusion pipeline before proceeding to the NeRF fine-tuning phase.
  • We find that fine-tuning the source NeRF with SWD loss can perform 3D style transfer well.
  • Our experimental results illustrate the rich capability of stylizing scenes with various text prompts.
2 Related Work
2.1 Implicit 3D Representation
NeRF, introduced by the seminal paper Mildenhall et al. (2020) became one of the most popular implicit 3D representation techniques due to several benefits. NeRF can render photo-realistic novel views with arbitrary resolution due to its continuous representation with a compact model compared to explicit representations such as polygon mesh or voxels. In our research, we use the "nerfacto" model implemented by Nerfstudio Tancik et al. (2023), which is a combination of modular features from multiple papers Wang et al. (2021); Barron et al. (2022); Müller et al. (2022); Martin-Brualla et al. (2021); Verbin et al. (2022) , designed to achieve a balance between speed and quality.
2.2 Style Transfer
2.2.1 2D Style Transfer.
Style transfer is a technique for blending images, a source image and a style image, to create another image that retains the first’s content but exhibits the second’s style. Since the introduction of the foundational style transfer algorithm proposed by Gatys et al. (2015), many follow-up works for 2D style transfer have been explored for further improvements such as faster optimization Johnson et al. (2016); Huang and Belongie (2017), zero-shot style-transfer Li et al. (2017), and photo-realism Luan et al. (2017).
2.2.2 3D Style Transfer.
Several recent 3D style transfer works have applied style transfer techniques using deep feature statistics Huang and Belongie (2017) to NeRF Liu et al. (2023); Wang et al. (2023); Zhang et al. (2022). Although these methods require a reference style image, text-based 3D editing techniques have also been proposed leveraging foundational 2D Text-to-Image (T2I) generative models. While Instruct 3D-to-3D Kamata et al. (2023) proposed using Score Distillation Sampling (SDS) loss Poole et al. (2022) for text guided NeRF stylization, Instruct-NeRF2NeRF Haque et al. (2023) and ViCA-NeRF Dong and Wang (2024) perform NeRF editing by optimizing the underlying scene with a process referred to as Iterative Dataset Update (Iterative DU), which gradually replaces the input images with edited images from InstructPix2Pix Haque et al. (2023), an image-conditioned instruction-based diffusion model, followed by an update of NeRF. Inspired by these methods, we also develop a 3D style transfer method for NeRF, supervised by images created by a diffusion pipeline but without Iterative DU.
2.3 Diffusion Models
Diffusion models Sohl-Dickstein et al. (2015); Song et al. (2020a); Dhariwal and Nichol (2021) are generative models that have gained significant attention for their ability to generate high-quality, diverse images. Inspired by classical non-equilibrium thermodynamics, they are trained to generate an image by reversing the diffusion process, progressively denoising noisy images towards meaningful ones. Diffusion models are commonly trained with classifier-free guidance Ho and Salimans (2022) to enable image generation conditioned on an input text.
2.3.1 Controlled Generations with Diffusion Models.
Leveraging the success of T2I diffusion models, recent research has expanded their application to controlled image generation and editing, notably in image-to-image (I2I) tasks Meng et al. (2021); Parmar et al. (2023); Kawar et al. (2023); Tumanyan et al. (2023); Mokady et al. (2023); Hertz et al. (2023a, 2022); Brooks et al. (2023). For example, SDEdit Meng et al. (2021) achieves this by first adding noise to a source image and then guiding the diffusion process toward an output based on a given prompt. ControlNet Zhang et al. (2023) was proposed as an add-on architecture for training T2I diffusion models with extra conditioning inputs such as depth, pose, edge maps, and more. Several recent techniques Hertz et al. (2023b); Sohn et al. (2024); Cheng et al. (2023) focus on generating style-aligned images. In our work, we use a depth-conditioned I2I pipeline with an attention-sharing mechanism similar to "StyleAligned" Hertz et al. (2023b) to create a set of multi-view images sharing a consistent style.
Figure 1: Overall Pipeline: Our method consists of distinct procedures. We first prepare a NeRF model of the source view images. Given the depth maps of the corresponding views (by either estimation or rendering by NeRF), we generate stylized multi-view images using a style-aligned diffusion model. Lastly, we fine-tune the source NeRF on the stylized images using the SWD loss.
3 Method
3.1 Preliminaries
3.1.1 Neural Radiance Fields.
NeRF Mildenhall et al. (2020) models a volumetric 3D scene as a continuous function by mapping a 3D coordinate $\mathbf{x}=(x,y,z)$ and a 2D viewing direction $\mathbf{d}=(\theta,\phi)$ to a color (RGB) $\mathbf{c}$ and a density ($\sigma$). This function $F_{\theta}:(\mathbf{x},\mathbf{d})\mapsto(\mathbf{c},\mathbf{\sigma})$ is often parameterized by a neural network combined with voxel grid structures or other encoding techniques Fridovich-Keil et al. (2022); Müller et al. (2022); Sun et al. (2022a, b) to accelerate performance. Given a NeRF model trained on a set of 2D images taken from various viewpoints of a target scene, the accumulated color $C(\mathbf{r})$ along an arbitrary camera ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ is calculated with the quadrature rule by volume rendering Max (1995):
$C(\mathbf{r})=\sum_{k=1}^{K}T_{k}(1-\exp{-\sigma_{k}\delta_{k}})\mathbf{c}_{k},T_{k}=\exp{\Bigl{(}-\sum_{j=1}^{k-1}\sigma_{j}\delta_{j}\Bigr{)}}$ (1)
where $\delta_{k}=t_{k+1}-t_{k}$ is the distance between sampled points on the ray and $T_{k}$ is the accumulated transmittance from origin $\mathbf{o}$ to the $k$-th sample.
3.1.2 Conditional Diffusion Models.
Recent T2I diffusion models Rombach et al. (2022); Podell et al. (2023) are built with a U-net architecture Ronneberger et al. (2015) integrated with convolutional layers and attention blocks Vaswani et al. (2017). Within the model, attention blocks play a crucial role in correlating text with relevant parts of the deep features during image generation. Our work uses an open-source latent diffusion model Podell et al. (2023), which includes a CLIP text encoder Radford et al. (2021) for text embedding. The cross-attention between contextual text embedding and the deep features of the denoising network is calculated as follows:
$\texttt{Attn}(Q,K,V)=\texttt{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}V\right)$ (2)
where $Q\in m\times d_{k},K\in m\times d_{k},V\in m\times d_{h}$ are projection matrices for a deep feature map $\phi\in\mathbb{R}^{m\times d_{h}}$. We may interpret the attention operation in equation 2 as values $V$, originating from conditional text, weighted by the correlation of queries $Q$, and the keys $K$. There are often $N_{h}$ multiple attention heads in each layer along the $d_{h}$ dimension to allow the model to attend to information from different subspaces in feature space jointly:
$\texttt{MultiHeadAttn}(Q,K,V)=\underset{j\in N_{h}}{\texttt{concat}}\Bigl{[}\texttt{Attn}(Q^{(j)},K^{(j)},V^{(j)})\Bigr{]}$ (3)
3.2 Style-NeRF2NeRF
Our method is a distinct two-step process. First, we prepare stylized images of corresponding source views using our style-aligned diffusion pipeline, and then refine the source NeRF model based on the generated views to acquire a style-transferred 3D scene.
3.2.1 Style-Aligned Image-to-Image Generation.
Given a set of source view images $\{I^{(i)}\}(i=1,\ldots,N)$, our first goal is to generate a corresponding set of stylized view images $I^{(i)}_{c}=U_{\theta}(I^{(i)},c)$ under a text condition $c$ with as much perceptual view consistencies among images where $U_{\theta}$ consists of a sampling process such as DDIM Song et al. (2020b).
Although T2I diffusion models can generate rich images with arbitrary text prompts, merely sharing the same prompt across different source views is insufficient to generate stylized images with a perceptually consistent style. To alleviate this problem, we apply a fully-shared-attention variant of a style-aligned image generation method proposed by Hertz et al. (2023b). Let $Q_{i},K_{i},V_{i}$ be the queries, keys, and values from a deep feature $\phi_{i}$ for view $I^{(i)}$, then we generate $n$ stylized views simultaneously using the following fully-shared-attention:
$\texttt{Attn}(Q_{i},K_{1\ldots n},V_{1\ldots n})\$ (4)
$K_{1\ldots n}=[K_{1},K_{2},\ldots K_{n}]^{T},V_{1\ldots n}=[V_{1},V_{2},\ldots V_{n}]^{T}$ (5)
Figure 4 illustrates an example of multi-view images generated with and without the fully-shared-attention mechanism.
3.2.2 Conditioning on Source Views.
To further strengthen perceptual consistencies across multi-view frames, we attach a depth-conditioned ControlNet Zhang et al. (2023) and optionally enable SDEdit Meng et al. (2021) for conditioning on the source view. As for the depth inputs, we may either render the corresponding depth maps from the source NeRF or use an off-the-shelf depth estimator model such as MiDaS Ranftl et al. (2020).
Given a set of translated multi-view images based on style text and their corresponding camera poses for training a source NeRF model, we may proceed to the NeRF refining stage described below.
3.2.3 NeRF Fine-Tuning.
Based on the perceptually view-consistent images $\{I^{(i)}_{c}\}$ created by the style-aligned image-to-image diffusion model, our next objective is to fine-tune the source NeRF scene to reflect the target style in a 3D consistent manner.
Although the stylized multi-view images are a good starting point for fine-tuning the source NeRF, we found that using a common RGB pixel loss is prone to over-fitting due to ambiguities in 3D geometry and color. Therefore, an alternative loss function that reflects the perceptual similarity is preferred for guiding the 3D style-transfer process. To meet our requirement, we employ the Sliced Wasserstein Distance loss (SWD loss) Heitz et al. (2021).
3.3 Sliced Wasserstein Distance Loss.
Feature statistics of pre-trained Convolutional Neural Networks (CNNs) such as VGG-19 Simonyan and Zisserman (2014) are known to be useful for representing a style of an image Gatys et al. (2015); Johnson et al. (2016); Huang and Belongie (2017); Li et al. (2017); Luan et al. (2017). In our study we employ the SWD loss originally proposed for texture synthesis Heitz et al. (2021) as the loss term to guide the style-transfer process for NeRF.
Let $F^{l}_{m}\in\mathbb{R}^{N_{l}}(m=1,\ldots,M_{l})$ denote the feature vector of the $l$-th convolutional layer at pixel $m$ where $M_{l}$ is the number of pixels and $N_{l}$ is the feature dimension size. Using the delta Dirac function, we may express the discrete probability density function $p^{l}(x)$ of the features for layer $l$ as below:
$p^{l}(x)=\frac{1}{M_{l}}\sum_{m=1}^{M_{l}}\delta_{F^{l}_{m}}(x)$ (6)
Using the feature distributions $p^{l},\hat{p}^{l}$ for image $I$ and its corresponding optimization target $\hat{I}$, the style loss is defined as a sum of SWD over the layers:
$\mathcal{L}_{style}=\sum_{l=1}^{L}\mathcal{L}_{SWD}(p^{l},\hat{p}^{l})$ (7)
where, $\mathcal{L}_{SWD}$ is the SWD term defined as the expectation over $1$-dimensional Wasserstein distances of features projected by random directions $V\in\mathcal{S}^{N_{l}-1}$ sampled from a unit hypersphere.
Using the projected scalar features $p^{l}_{V}=\{\langle F^{l}_{m},V\rangle\},\forall m$, where $\langle$,$\rangle$ , ⟩ denotes a dot product, one may obtain $\mathcal{L}_{SWD}$ as the following where the $1$-dimentional $2$-Wasserstein distance $\mathcal{L}_{SW1D}$ is trivially calculated in a closed form by taking the element-wise $L^{2}$ distances between sorted scalars in $p^{l}_{V}$ and $\hat{p}^{l}_{V}$. An illustration of a projected 1D Wasserstein distance is shown in figure 2.
$\mathcal{L}_{SWD}=\sum_{l=1}^{L}\mathbb{E}_{V}[\mathcal{L}_{SW1D}(p^{l}_{V},\hat{p}^{l}_{V})]$ (8)
$\mathcal{L}_{SW1D}(p^{l}_{V},\hat{p}^{l}_{V})=\frac{1}{|p^{l}_{V}|}\lVert\texttt{sort}(p^{l}_{V})-\texttt{sort}(\hat{p}^{l}_{V})\rVert^{2}$ (9)
Expectation over random projections $V$ provides a good approximation in practice and an optimized distribution is proven to converge to the target distribution. SWD is known to capture the complete target distribution Pitie et al. (2005) as described below:
$\mathcal{L}_{SW}(I,\hat{I})=0\implies p^{l}=\hat{p}^{l},\forall l\in{1,\ldots,L}$ (10)
The calculation of SWD scales in $\mathcal{O}(M\log M)$ for an $M$-dimensional distribution, making it suitable for machine learning applications with gradient descent algorithms.
Figure 2: Sliced Wasserstein Distance: $p$ and $\hat{p}$ are projected onto a random unit direction $V$ (left). The $1$-dimensional Wasserstein distance can be calculated by taking the $L^{2}$ difference between the sorted projections $p$ and $\hat{p}$ (right). Expectation over random $V$ vectors is a practical approximation of the $N$-dimensional Wasserstein distance.
3.4 Style Blending.
Given two different stylized views $I_{1},I_{2}$ and their corresponding feature distributions $\{p_{1}^{l},p_{2}^{l}\}$, one may obtain a style-blended scene by refining the source NeRF model towards the Wasserstein barycenter where $t\in[0,1]$ is the blending weight between the two styles:
$\mathcal{L}_{style}(I_{1},I_{2},\hat{I})=\sum_{l=1}^{L}\left(t\mathcal{L}_{SWD}(p_{1}^{l},\hat{p}^{l})+(1-t)\mathcal{L}_{SWD}(p_{2}^{l},\hat{p}^{l})\right)$ (11)
An example of style blending is shown in figure 3.
Figure 3: Style Interpolation: An example of style blending using the Wasserstein barycenter between two different style prompts "A person like Marilyn Monroe, pop art style" and "A person like Steve Jobs".
3.5 Implementation Details.
We employ Stable Diffusion XL Podell et al. (2023) as a backbone for the style-aligned image-to-image diffusion pipeline. For NeRF representation, we use the "nerfacto" model implemented in NeRFStudio Tancik et al. (2023). Due to memory constraints, we generate up to $18$ views simultaneously with a fixed seed across all $N$ source views. In our experiments, we generated the target images with $50$ denoising steps using a range of classifier-free guidance weights (mostly between 5 and 30) depending on the scene or the style text. While we use depth maps rendered by NeRF for relatively compact and forward-facing scenes, we opt for depth estimations from the MiDaS model Ranftl et al. (2020) for large-scale outdoor scenes.
As the image editing is performed before NeRF training, our method allows users to test with different text prompts and parameters (e.g., text guidance scale, SDEdit strength) beforehand. Additionally, our straight-forward NeRF training without Iterative DU Haque et al. (2023) or score distillation sampling (SDS) loss Poole et al. (2022) allows the training process to run with less GPU memory as editing by a diffusion model is not necessary during NeRF updates. We also verify the importance of style-alignment in the ablation study. Please refer to the supplementary material for more implementation details. The overall pipeline of our method is shown in Figure 1.
Figure 4: Effect of Style-Alignment: An example of source view conversion applied to "Bear" scene using a text prompt "A bear, impressionism painting style" with and without shared-attention mechanism within the diffusion pipeline. We find that a fully-shared-attention variant of the style-aligned diffusion model Hertz et al. (2023b) greatly improves perceptual consistencies among generated views.
4 Results
We run our experiments on several real-world scenes, including the Instruct-NeRF2NeRF Haque et al. (2023) dataset captured by a smartphone and a mirrorless camera with camera poses extracted by COLMAP Schönberger and Frahm (2016) and PolyCam pol . The dataset contains large-scale 360 scenes, objects, and forward-facing human portraits. We show qualitative results and comparisons against several variants to verify the effectiveness of our method design with CLIP Text-Image Direction Similarity (CLIP-TIDS), a quantitative metric originally introduced in StyleGAN-Nada Gal et al. (2022) and CLIP Directional Consistency (CLIP-DC), a score proposed by Instruct-NeRF2NeRF Haque et al. (2023) that aims to measure the view-consistency between adjacent frames. In addition to the above, we also present comparison results against recent NeRF-based 3D editing methods, Instruct-NeRF2NeRF Haque et al. (2023) and ViCA-NeRF Dong and Wang (2024). We encourage our readers to see the results in the supplementary video.
Figure 5: Baseline Comparisons: We compare our method against several variants. The images show an example comparison of the "Bear" scene edited with a text description "A bear, impressionism painting style" with a text guidance scale of $7.5$. Note that (b), (c), (e), and (f) are all novel view renders from NeRF. We encourage our readers to check the results in the video.
Figure 6: Qualitative Results: We show novel view rendering examples of real-world scenes stylized or edited with text descriptions specifying certain artistic styles or environmental changes such as weather conditions.
Figure 7: Method Comparison: A comparison of NeRF stylization methods. While we used a text guidance scale of between $15$ to $25$ for our results, it is controllable via text prompts concerning subjective preferences. Note that all images are novel view renders from NeRF.
4.1 Qualitative Evaluation
Qualitative results are shown in Figure 6. Our method is capable of performing artistic style transfer under various style prompts without hallucinations. We recommend watching the supplementary video to confirm that the stylized scenes are sufficiently view-consistent.
4.2 Ablations
We verify the effectiveness of our method by comparing it against the following variants. An illustration of the comparison results is shown in Figure 5.
  • No Style-Alignment: To examine the importance of preparing perceptually view-consistent stylized images prior to the training process of a source NeRF model, we turn off the full-attention-sharing. Due to the view-inconsistencies in stylized images (See also middle row in figure 4), fine-tuning NeRF on such images will result in an unpredictable mixture of styles.
  • Style-Alignment Train-from-Scratch: In this naive variant, we train a NeRF from scratch using the images generated with our style-aligned diffusion pipeline. Without pre-training of the underlying scene, 3D style transfer produces floating artifacts and shape inconsistencies due to ambiguities in geometry and color of stylized training images.
  • Style-Alignment w/ RGB Loss: This variant trains the source NeRF with $L^{1}$ pixel RGB loss instead of SWD loss. As perceputal view-consistency or similar style does not guarantee physically consistent geometry and color across different views, training with RGB loss tends to diverge to a blurry scene. RGB loss is prone to over-fitting, whereas SWD is a more valid choice for effectively learning the perceptual similarity from style-aligned training images.
4.2.1 Quantitative Evaluation.
We quantitatively measure our method against the variants using CLIP-TIDS and CLIP-DC with a fixed text guidance scale of 15. The results are shown in table 1.
Table 1: Quantitative Evaluation We show both CLIP-TIDS and CLIP-DC results. The values are the average of the novel view render frames over two scenes using five prompts.
No Style-Align Train from Scratch Style-Align w/ RGB Loss Ours
CLIP-TIDS $0.125$ $0.073$ $0.237$ $\mathbf{0.162}$
CLIP-DC 0.932 0.917 0.917 0.928
4.3 Method Comparison
We compare our method against Instruct-NeRF2NeRF Haque et al. (2023) and ViCA-NeRF Dong and Wang (2024) on four scenes including two large-scale outdoor scenes, a 360 object scene (Bear) and a forward-facing scene (a human portrait) using three different text prompts for each scene. Our method exhibits competitive style transfer results whereas previous methods occasionally suffer from hallucination effects (e.g. The Janus problem etc…) caused by the underlying diffusion model. As the generation of images and NeRF refinement is a separate process in our method, it is possible to filter out and recreate any images that could have undesired impact on the NeRF fine-tuning. Visual results are given in figure 7.
4.3.1 User Study
For each scene, participants were shown a combination of stylized views rendered by different methods in random order, and were asked to select a single view that most likely adhere to the provided stylization text prompt. In our user study, we collected feedbacks from 33 individuals resulting in a total of 396 votes. The overall percentage of the selected preferred method is shown in table 2, indicating that our method can perform competitive artistic style transfer without hallucinations.
4.3.2 Quantitative Comparison
As style transfer is inherently a subjective task, we think that qualitative evaluation by the user is the most important. Nevertheless, we additionally provide quantitative comparison results using CLIP-TIDS and CLIP-DC. Results are included in table 2.
Table 2: Method Comparison Results The metrics are the average of novel view renders over four scenes with each using three prompts. (Average of $4\times 3=12$ style transfer results.) Our method shows the best values for CLIP-TIDS, CLIP-DC, and user preference.
CLIP-TIDS CLIP-DC User Preference
Instruct-NeRF2NeRF $0.081$ $0871$ $15.3\%$
ViCA-NeRF $0.061$ $0.914$ $17.6\%$
Ours 0.084 0.923 $\mathbf{67.1\%}$
5 Limitations and Future Work
While our method may apply artistic style transfer to various scenes, including large-scale outdoor environments, there are several limitations that we’d like to consider. Thin structures such as plants and trees in the background or delicate texture patterns are challenging to reconstruct due to ambiguities in the stylized multi-view images. For the same reason, our method will struggle to learn fine details if there is too much variation in the training images (e.g. different people or objects in the background, random patterns of clouds in the sky). As the style-aligned diffusion pipeline is conditioned on depth maps, significant editing of geometry is also difficult.
We think our approach is applicable to other types of 3D representations such as 3D Gaussian Splatting Kerbl et al. (2023) and extendable to more features such as scene relighting and deformation, which are exciting directions for further exploration.
6 Conclusion
We propose a novel 3D style-transfer method for NeRF representation leveraging a style-aligned generative diffusion pipeline. By guiding the training process with Sliced Wasserstein Distance or SWD loss, the source 3D scene, pre-trained as a NeRF model, is effectively translated into a stylized 3D scene. The method is a relatively straightforward two-step process, allowing the creators to visually search and refine their style concepts by testing various text prompts and guidance scales before fine-tuning the source NeRF model. Our proposed method shows competitive 3D style transfer results compared to previous methods and can blend styles by optimizing the source 3D scene towards the Wasserstein barycenter.
7 Acknowledgements
This work was partially supported by JST Moonshot R&D Grant Number JPMJPS2011, CREST Grant Number JPMJCR2015, and Basic Research Grant (Super AI) of Institute for AI and Beyond of the University of Tokyo. We want to thank Instruct-NeRF2NeRF Haque et al. (2023) authors for sharing their dataset, and Yuki Kato, Shinji Terakawa, and Yoshiaki Tahara for assisting with data capturing.
Appendix A Additional Implementation Details
We pre-train the "nerfacto" model implemented in Nerfstudio Tancik et al. [2023] for 60,000 iterations and then fine-tune for 15,000 iterations. We use the default "nerfacto" losses; RGB pixel loss $\mathcal{L}_{rgb}$, distortion loss $\mathcal{L}_{dist}$ Barron et al. [2022], interlevel loss $\mathcal{L}_{inter}$ Barron et al. [2022], orientation loss $\mathcal{L}_{orien}$ Verbin et al. [2022], and predicted normal loss $\mathcal{L}_{normal}$ Verbin et al. [2022] for pre-training (equation 12). During fine-tuning, we disable the RGB pixel loss $\mathcal{L}_{rgb}$, orientation loss $\mathcal{L}_{orien}$, and predicted normal loss $\mathcal{L}_{normal}$ but add the Sliced Wasserstein Distance (SWD) loss Heitz et al. [2021] (equation 13). The total loss function for each phase is as follows:
$\mathcal{L}_{pre}=\mathcal{L}_{rgb}+\lambda_{1}\mathcal{L}_{dist}+\lambda_{2}\mathcal{L}_{inter}+\lambda_{3}\mathcal{L}_{orien}+\lambda_{4}\mathcal{L}_{normal}$ (12)
$\mathcal{L}_{fine}=\mathcal{L}_{swd}+\lambda_{1}\mathcal{L}_{dist}+\lambda_{2}\mathcal{L}_{inter}$ (13)
where we use the default hyper-parameters $\lambda_{1}=0.002,\lambda_{2}=1.0,\lambda_{3}=0.0001,\lambda_{4}=0.001$ for most cases. A greater value for the distortion loss may work better if floating artifacts remain in the scene. We list a brief description of the "nerfacto" losses.
  • Distortion Loss: The loss encourages the density along a ray to become compact, aiming to prevent floaters and background collapse. It was proposed in Mip-NeRF 360 Barron et al. [2022].
  • Interlevel Loss: The loss allows the histograms of the point sampling proposal network and NeRF network to become more consistent. It was also proposed in Mip-NeRF 360 Barron et al. [2022]
  • Orientation Loss: The loss aims to prevent "foggy" surfaces by penalizing visible samples with predicted normals facing the ray direction. It was introduced in Ref-NeRF Verbin et al. [2022].
  • Predicted Normal Loss: The loss enforces the predicted normals to be consistent with density gradient normals. It is often used in conjunction with the orientation loss.
For detailed definitions of the "nerfacto" losses, please see Mip-NeRF Barron et al. [2022] and Ref-NeRF Verbin et al. [2022].
The SWD loss $\mathcal{L}_{swd}$ is applied to $64\times 64$ patches sampled during NeRF fine-tuning. Although we found that $64\times 64$ empirically works sufficiently, one may change the patch size accordingly. We run all our experiments with Python 3.10 and CUDA 11.8 on a single NVIDIA H100. Although we optionally enable SDEdit Meng et al. [2021] in our style-aligned diffusion pipeline, we recognized that depth maps are often enough for conditioning on the original views. In such cases, we use $1.0$ as the strength for SDEdit.
Appendix B CLIP Text-Image Direction Similarity
CLIP Text-Image Direction Similarity (CLIP TIDS) Gal et al. [2022] is a metric for evaluating how well the change in the stylized image is aligned with the user-provided text prompt. Given the CLIP image and text encoder $E_{I},E_{T}$, CLIP TIDS between the source and the stylized image $I_{source},I_{style}$ is calculated as:
$\Delta I=E_{I}(I_{style})-E_{I}(I_{source}),\Delta T=E_{T}(t_{style})-E_{T}(t_{source})$ (14)
[…]
where $t_{source},t_{style}$ are text descriptions describing the images (e.g. $t_{source}=$"A photo of a person", $t_{style}=$"A person like Vincent Van Gogh").
Appendix C Effect of Text Guidance Scale
In figure 8, we show an example of NeRF renderings using the same style prompt "A person like Vincent Van Gogh" but different text guidance scales. We may see that style strength is controllable while keeping the original content structure grounded on the original image. We also show CLIP TIDS for each text guidance scale in table 3. A stronger text guidance scale will lead to higher CLIP TIDS.
Table 3: CLIP-TIDS Comparison CLIP-TIDS values over test renders for each text guidance scale are shown. We can verify that a stronger text guidance scale results in higher CLIP TIDS values.
Text Guidance Scale $s=7.5$ $s=15.0$ $s=22.5$ $s=30.0$ $s=37.5$
CLIP TIDS $0.0693$ $0.1459$ $0.1520$ $0.1594$ 0.1718
Figure 8: Effect of Text Guidance Scale: We show some NeRF rendering results for the prompt "A person like Vincent Van Gogh" using various text guidance scales ($s=7.5,15.0,22.5,30.0,37.5$).
Appendix D Limitations
While our method can effectively perform overall style transfer to 3D scenes, it is still difficult to reconstruct a detailed structure of fluctuating objects within the stylized multi-view images. In figure 9, for example, (c) the clouds in the fine-tuned NeRF scene show a fractal-like pattern, which is different from (b) the clouds illustrated in the stylized image. This phenomenon is due to the ambiguities of cloud positions or shapes appearing in the stylized images generated by the style-aligned diffusion pipeline Hertz et al. [2023b]. We leave the development of a more robust content structure-preserving style transfer technique as future work.
Figure 9: Limitations: Due to remaining ambiguities in the stylized multi-view images, fluctuating objects such as clouds may lose their detailed shape in the fine-tuned NeRF renderings. We wish to improve on this in our future work.